# GASP: Gaussian Splatting for Physic-Based Simulations

### v1

@@ Abstract
Physical simulation is paramount for modeling and utilization of 3D scenes in various real-world applications. However, its integration with state-of-the-art 3D scene rendering techniques such as Gaussian Splatting (GS) remains challenging. Existing models use additional meshing mechanisms, including triangle or tetrahedron meshing, marching cubes, or cage meshes. As an alternative, we can modify the physically grounded Newtonian dynamics to align with 3D Gaussian components. Current models take the first-order approximation of a deformation map, which locally approximates the dynamics by linear transformations. In contrast, our Gaussian Splatting for Physics-Based Simulations (GASP) model uses such a map (without any modifications) and flat Gaussian distributions, which are parameterized by three points (mesh faces). Subsequently, each 3D point (mesh face node) is treated as a discrete entity within a 3D space. Consequently, the problem of modeling Gaussian components is reduced to working with 3D points. Additionally, the information on mesh faces can be used to incorporate further properties into the physical model, facilitating the use of triangles. Resulting solution can be integrated into any physical engine that can be treated as a black box. As demonstrated in our studies, the proposed model exhibits superior performance on a diverse range of benchmark datasets designed for 3D object rendering. Our project page is at: https://waczjoan.github.io/GASP/.
@@ 1 Introduction
Figure 1: GASP integrates GS with a physical engine to generate realistic simulations. Essentially, GASP utilizes set of triangle meshes to parameterize Gaussian components. This approach reduces modifying Gaussian components to 3D point clouds (consisting of vertices of triangles) processing, enabling efficient and rapid rendering. GASP can model the interaction of objects and work with dynamic scenes.
@@
The recent development of Gaussian Splatting (GS) [13] has greatly influenced the field of computer graphics, enabling the modeling of three-dimensional (3D) scenes from images with annotated camera positions. The GS model represents 3D scenes as a set of 3D Gaussian distributions, which allows for rapid training and rendering. Furthermore, it is relatively straightforward to modify the original GS approach to enable editing [22, 4, 8] and to facilitate the creation of dynamic scenes [21, 11, 14].
Figure 2: GASP produces physical simulations in flat GS. It works with static GS models like GaMeS [22] or dynamic ones like D-MiSo [21]. (In general, GASP can be applied to any GS-based model which uses flat Gaussians.) First, we take the flat Gaussian and convert it to the GaMeS format. Each flat Gaussian component is converted to three points (triangle face). Then, the physical engine is run on a point cloud to obtain trajectories for our 3D model. Finally, we convert points into Gaussian components by reverse GaMeS parametrization.
At the same time, one of the most significant challenges with the GS framework is the incorporation of physics into scenes represented using 3D Gaussian components. It is noteworthy that using a physical engine would pave the way to the implementation of GS in spatial, immersive interfaces such as virtual reality (VR) [12]. Presently, the existing models employ supplementary meshing techniques, including triangle or tetrahedron meshing [2], marching cubes [4], or cage meshes [12]. This approach generates classical mesh-based representations, which are then controlled by a physical engine. Consequently, Gaussian components are adjusted in accordance with mesh-based modifications. While such methods produce satisfactory renderings, they require the implementation of an additional meshing and rendering strategy. Furthermore, modeling tears and cracks represents a challenge, as it is not straightforward to control the behaviors of the mesh when the object is divided. An alternative methodology for addressing the aforementioned issue is to utilize the vanilla GS approach and adapt Newtonian dynamics to accommodate 3D Gaussian components. PhysGaussian [24] employs the first-order approximation of a deformation map. By utilizing such a method, we locally approximate the dynamics through linear transformations. This solution is effective, however, it is contingent upon access to a physical engine in order to apply the modifications and relies on a network approximation.
@@
To address these limitations, we propose the Gaussian Splatting for Physic-Based Simulations (GASP) [^1] The code is avaliable at https://github.com/waczjoan/GASP model (see Fig. 1), which incorporates flat Gaussians [22, 21] with the Material Point Method (MPM) [5] without additional modification (see Fig. 2). Flat Gaussian distributions can be used in GS by applying only minimal changes, such as imposing a zero value on one of the Gaussian components. Such techniques are frequently employed in a multitude of approaches [22, 4, 7], and they facilitate straightforward modifications to the final model. The training and rendering times, as well as the quality of the renderings produced by flat Gaussians, are comparable to those of GS. GASP approach exemplifies the implementation of Newtonian dynamics within the GaMeS framework, thereby illustrating the capacity to integrate these methodologies without the need for external modifications to the physical engine or Gaussian components. More precisely, the model extracts points from GS and applies the Material MPM to control such points. During the simulation process, the Gaussian components associated with each triple of points are recalculated. This process is notably rapid and does not increase rendering costs. At the same time, treating each point as a discrete entity results in the emergence of artifacts. Consequently, we propose a simple modification that utilizes triangles instead of completely disparate points. We use additional rescaling when a triangle undergoes a significant change in size. In practice, GASP can produce high-quality simulations in flat GS, including static and dynamic scenes (see Fig. 3) as well as the interaction of objects (see Fig. 4). This is evidenced by the results of experiments conducted on a range of benchmarks with synthetic (see Fig. 5 ) and real-world data sets (see Fig. 6 ) .
@@
    We propose GASP model that incorporates physical properties into a 3D scene representation using GS on both static and dynamic scenes.
@@
    GASP operates in a direct manner on specific points, thus obviating the necessity for alterations to the physical engine regarded as a black box.
@@
    GASP does not employ supplementary meshing strategies and instead operates directly with Gaussian distributions.
@@ 2 Related Work
The GS representation of 3D scenes is well suited for editing [22, 4, 7] and dynamic scenes modeling [10, 23, 21]. However, the more challenging task is its integration with a physical engine that remains difficult and underexplored.
Figure 3: In this experiment, we added simulated gravity and lowered a rubber teddy bear. As you can see, when the object hits the ground, it changes according to the properties given by the physics engine.
@@
Most of the existing methods use additional strategic tactics to obtain representation well-suited for physical engines. Gaussian Splashing (GSP) [2] is a unified framework that combines 3D Gaussian Splinting (3DGS) and position-based dynamics. During the preprocessing stage, the foreground objects are isolated and reconstructed as a mesh representation. Such a response is then combined with position-based dynamics. Alternatively, we can use cage-based representation [12]. VR-GS [12] combine 3DGS and eXtended Position-based Dynamics (XPBD) [16]. The model starts with scene reconstruction by GS and tetrahedralization. Consequently, we obtain a cage-based representation combined with a physical engine.
@@
All the above models work with an additional meshing strategy. In contrast, PhysGaussian [24] operates directly on Gaussian distributions. The authors of [24] suggest approximating a deformation map using the first-order derivative. This yields a linear transformation capable of modifying Gaussian components. Although this model produces relatively good results, it requires adjustments to the physical engine. In contrast, we propose GASP model which uses a physical engine as a black-box without any modification.
@@
A separate area in the physics modeling involves using diffusion models to generate object dynamics. PhysDreamer [25] introduces a method that provides static 3D objects with interactive dynamics by utilizing object dynamics priors acquired from video generation models. Physics3D [15] learns various physical characteristics of 3D objects using a video diffusion model. DreamPhysics [9] infers physical attributes of 3DGS based on video diffusion priors.
3 GASP Model Overview
Figure 4: GASP reduces the modification of the Gaussian component to 3D point cloud motion. Therefore, we can work with several elements at the same time. In this experiment, we simulate a static mug and a teddy bear made of sand.
Figure 5: We add small elements to large scenes that interact with them. At the top, we add a rubber duck that falls onto the table. At the bottom, we add the sand-based object. Note that GASP is able to model interaction between 3D objects.
Figure 6: In large scenes, we can select part of the object and add physical properties. In this experiment, we simulate the effect of wind to modify the movement of the bonsai tree.
Figure 7: Blender provides a wide range of physical properties and force sources, including gravity and wind, to create realistic and dynamic animations. In this experiment, we present examples of physical animation in object scenes, showing the versatility of animating both small objects, such as a flower on a table, and large objects, such as a fox’s head dominating the scene.
@@ 3 GASP Model Overview
Here, we present a description of the GASP model. It begins with a general GS model, followed by an examination of the modifications made by GaMeS [22] to the original GS model, which facilitates straightforward adaptations. Subsequently, the classical MPM [5] is outlined, and finally, our own approach combining both GS and MPM is detailed.
Gaussian Splatting
The GS technique [13] generates 3D scenes using a collection of three-dimensional Gaussians:
$\mathcal{G}=\{(\mathcal{N}({\bf m}_{i},\Sigma_{i}),\sigma_{i},c_{i})\}_{i=1}^{p},$ (1)
defined by their mean (position) ${\bf m}_{i}$, covariance matrix $\Sigma_{i}$, opacity $\sigma_{i}$, and color $c_{i}$, which is represented using spherical harmonics (SH) [3]. The performance of the GS algorithm is mainly due to its rendering technique, which projects the Gaussian elements into two-dimensional space. Throughout training, all parameters are optimized according to the mean square error (MSE) cost function. Since this training often leads to local minima, GS may utilize additional training methods that include creating, removing, and repositioning components based on the proposed heuristic [13]. This strategy is both fast and efficient.
@@
In the classical GS model, each element is characterized by a collection of parameters, including a covariance matrix $\Sigma$ which is factored as:
$\Sigma=RSSR^{T},$ (2)
where $R$ is the rotation matrix and $S$ is a diagonal matrix containing the scaling parameters. In GaMeS [22], on the other hand, flat Guassians are used:
$(\mathcal{N}({\bf m},R,S),\sigma,c)$ (3)
where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$. Thus, Gaussian components can be visualized through a triangle-face mesh by utilizing the vertices for their parameterization. We refer to this mapping as $\mathcal{T}(\cdot)$. In practice, this parameterization results in a set of triangles, which is referred to as Triangle Soup [18].
To outline the GaMeS parameterization, consider a single Gaussian component $\mathcal{N}({\bf m},R,S)$. Then its face representation $\mathcal{N}(V)$ is based on a triangle:
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ (4)
with the vertices defined as:
${\bf v}_{1}={\bf m},\;\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\;\textrm{and}\;\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ (5)
Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
@@
Here, $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process [1]. Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as:
$\hat{\Sigma}=\hat{R}\hat{S}\hat{S}\hat{R}^{T}.$ (10)
@@
which is used to measure local rotation and strain. The deformation function $\phi$ is subject to changes in accordance with the principles of conservation of mass, conservation of momentum, and the elasto-plastic constitutive relation:
@@
Figure 8: GASP approach operates directly on Gaussian representations without altering the physics-based deformation network. Initially, the 3D objects are depicted as flat Gaussians. GaMeS then converts these Gaussians into triplets (mesh faces). The simulation engine transforms these points following the MPM model. Subsequently, the new points are reassembled into triangles and then into Gaussians. This process is extremely fast, yielding realistic renderings.
@@
This subsection presents a description of the GASP model. Since the model employs the GaMeS representation of a 3D scene, we postulate that the input comprises a collection of Gaussians:
@@
with means ${\bf m}_{i}$, scaling matrices $S_{i}$, and rotation matrices $R_{i}$. Furthermore, we assume the existence of a deformation map $\phi(X,t)$, which models the behavior of 3D objects. In order to obtain $\phi(X,t)$, it is necessary to have a point cloud that describes 3D objects. As our method employs the GaMeS reparametrization of GS, we are able to transform the mean Gaussian parameters and the covariance matrix into a mesh (triangle soup). In such an instance, we may consider the system $(\mathcal{V},\mathcal{E})$, where:
@@
denotes a collection of vertices, and $\mathcal{E}$ represents the edges obtained via GaMeS. Our physical-based deformation network operates on each point of $V_{i}$ individually. Subsequently, the vertices of the triangle soup can be used as input for the MPM algorithm in conjunction with tools such as the Blender or Taichi libraries (for further details, see the experimental section). It is essential to note that our solution does not entail any modifications to the MPM, in contrast to the Taylor approximation employed in [24]. Consequently, in each time step, the following is obtained:
@@
The mesh in time $t$ is then defined by the collection of vertices and edges $(\mathcal{V}_{t},\mathcal{E}_{t})$. In contrast, the reverse parametrization provided by GaMeS can be employed to obtain Gaussian Splatting:
@@
It is important to note that processing each vertex of a triangle soup individually is a highly rapid process that treats the physical engine as an opaque black box. The schema of such processes is presented in Fig. 8. Our method is compatible with any physics engine that operates on three-dimensional points. However, the processing of all points separately may yield artifacts. When all points are treated independently during simulation, some elements exhibit considerable fluctuations in velocity. This phenomenon leads to the formation of extensive triangles (mesh faces), which in turn give rise to prominent Gaussian components, as illustrated in Fig. 9. To mitigate this effect, our approach employs a straightforward yet highly effective technique (correction algorithm). Specifically, we monitor the distance between vertices of the face representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ of a Gaussian component $\mathcal{N}({\bf m},R,S)$. It should be noticed that in the case of GaMeS transformation, the first vertex is equal to the center of the distribution (${\bf v}_{1}={\bf m}$), while the two subsequent vertices are located on the main components of the ellipses that describe the distribution (${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$). Consequently, if the distance between the transformed mesh nodes is considerably greater than that of the original nodes, i.e.:
@@
where $\alpha$ is a hyperparameter, we automatically assert that the value of $s_{i}$ should be identical to that of the original value times $\alpha$. This is illustrated in Fig. 10. It should be noted that throughout the full simulation, both the mesh and GS are accessible, enabling real-time rendering.
Figure 9: Comparison of physical animation models, illustrating the effects of applying an algorithm for correcting points used in simulation. The left side demonstrates a simulation without the correction algorithm, while the right side shows the enhanced accuracy and stability achieved using Taichi elements with the correction algorithm applied.
Figure 10: GASP works directly on points cloud extracted from flat Gaussians. In practice, physical engine moves all particles independently, which can produce artifacts. Therefore, we use a correction algorithm to control the distance of points from each other. If some points are moved extremely far, we automatically correct such behaviors by clipping parameters $S$ in GS.
@@
GASP is a natural combination of a physical engine dedicated to points and Gaussian Splatting with flat Gaussians. The reparametrization in GaMeS can be applied to a multitude of disparate methodologies, with a particular focus on flat Gaussian methods [4, 7]. Moreover, it can be employed for dynamic Gaussian Splatting. In [21], D-MiSo model is introduced, which uses flat Gaussians for the modeling of dynamic scenes. Such representations can be utilized in conjunction with the aforementioned approach. In the D-MiSo framework, two distinct types of Gaussians and two deformable networks are employed. In the process of inference, only sub-Gaussians are employed, which are modified by deformable function composition $\psi(v,s)$, i.e.,
@@
where $s$ represents a time value derived from the dynamics observed in the scenes. In our solution, we can incorporate a deformation map $\phi(X,t)$, which originates from the physics simulator. Consequently, GASP for the dynamic scene is given as follows:

### v3

@@ Abstract
Physics simulation is paramount for modeling and utilizing 3D scenes in various real-world applications. However, integrating with state-of-the-art 3D scene rendering techniques such as Gaussian Splatting (GS) remains challenging. Existing models use additional meshing mechanisms, including triangle or tetrahedron meshing, marching cubes, or cage meshes. Alternatively, we can modify the physics-grounded Newtonian dynamics to align with 3D Gaussian components. Current models take the first-order approximation of a deformation map, which locally approximates the dynamics by linear transformations. In contrast, our GS for Physics-Based Simulations (GASP) pipeline uses parametrized flat Gaussian distributions. Consequently, the problem of modeling Gaussian components using the physics engine is reduced to working with 3D points. In our work, we present additional rules for manipulating Gaussians, demonstrating how to adapt the pipeline to incorporate meshes, control Gaussian sizes during simulations, and enhance simulation efficiency. This is achieved through the Gaussian grouping strategy, which implements hierarchical structuring and enables simulations to be performed exclusively on selected Gaussians. The resulting solution can be integrated into any physics engine that can be treated as a black box. As demonstrated in our studies, the proposed pipeline exhibits superior performance on a diverse range of benchmark datasets designed for 3D object rendering. The project webpage, which includes additional visualizations, can be found at https://waczjoan.github.io/GASP.
@@
Figure 1: GASP integrates Gaussian Splatting (GS) with a physics engine to generate realistic simulations. Essentially, GASP utilizes a set of triangles to parameterize Gaussian components. This approach reduces modifying Gaussian components to 3D point clouds (consisting of vertices of triangles) processing, enabling efficient and rapid rendering. GASP can model the interaction of objects and work with dynamic scenes.
@@ 1 Introduction
The recent development of Gaussian Splatting (GS) [13] has greatly influenced the field of computer graphics, enabling the modeling of three-dimensional (3D) scenes from images with annotated camera positions. The GS model represents 3D scenes as a set of 3D Gaussian distributions, allowing rapid training and rendering. Furthermore, it is relatively straightforward to modify the original GS approach to enable editing [5, 9, 22] and to facilitate the creation of dynamic scenes [11, 14, 21].
At the same time, one of the most significant challenges with the GS framework is the incorporation of physics into scenes represented using 3D Gaussian components. Existing models employ additional meshing techniques, including triangle or tetrahedron meshing [4], marching cubes [5], or cage meshes [12]. This approach generates classical mesh-based representations, which are then controlled by a physics engine. Consequently, Gaussian components are adjusted following mesh-based modifications. Although such methods produce satisfactory renderings, they require the implementation of an additional meshing and rendering strategy. Extracting meshes from 3DGS remains a subject of active research, with promising approaches such as SuGaR or 2DGS [8]. However, obtaining smooth, high-quality meshes for arbitrary objects or scenes is still a challenging task, and artifacts affecting the simulation, such as mesh gaps, may occur.
Alternatively, PhysGaussian [24] utilizes the vanilla GS approach and adapts Newtonian dynamics to accommodate 3D Gaussian components. PhysGaussian employs the first-order approximation of a deformation map. Using such a method, it can locally approximate the dynamics through linear transformations. This solution is effective, however, it is contingent upon access to a physics engine in order to apply the modifications and relies on a network approximation. Due to that, this method is not suitable for use in other popular 3D graphics software such as Blender.
To address these limitations, we propose the Gaussian Splatting for Physics-Based Simulations (GASP) pipeline (see Fig. 1), which incorporates flat Gaussians [21, 22] with black box physics engine without special modifications to GS-based model which makes it more flexible compared to other methods. In this work, we present simulations created using the Genesis[2], Blender software and the Taichi Elements library (see Fig. 2). Additionally, to the best of our knowledge, this is the first demonstration of the integration of a GS-based model with the Genesis engine. We also demonstrate that straightforward enhancements, such as implementing the Gaussian hierarchy, can seamlessly integrate into our method, significantly accelerating the simulation process.
Figure 2: The GASP pipeline is versatile and compatible with various engines, requiring no significant modifications for integration. We demonstrate its capabilities with simulations performed using the Genesis engine, Blender software, and Taichi Elements.
@@
Flat Gaussian distributions can be smoothly employed in GS by modifying Gaussian’s scaling component, so it has a zero value on one axis. This representation of GS enables straightforward parametrization and can be easily converted into faces compatible with traditional 3D graphic software. Such techniques have been previously used in several approaches [5, 8, 22]. The training and rendering times, as well as the quality of the renderings produced by flat Gaussians, are comparable to those of GS. GASP approach exemplifies the implementation of Newtonian dynamics within the GaMeS framework, thereby illustrating the capacity to integrate these methodologies without needing external modifications to the physics engine or Gaussian components. More precisely, the model extracts points from GS and applies the Material Point Method to control such points. During the simulation process, the Gaussian components associated with each triple of points are recalculated. This process is notably rapid and does not increase rendering costs. At the same time, treating each point as a discrete entity results in the emergence of artifacts. Consequently, we propose a simple modification that utilizes triangles instead of completely disparate points. We use additional rescaling when a triangle undergoes a significant change in size. In practice, GASP can produce high-quality simulations in flat GS, including static and dynamic scenes as well as the interaction of objects, see Fig. 1. This is evidenced by the results of experiments conducted on a range of benchmarks with synthetic and real-world data sets.
@@
    We propose GASP pipeline that incorporates physical properties into a 3D scene representation using GS-based models on both static and dynamic scenes.
@@
    GASP operates directly on specific points, thus obviating the necessity for alterations to the physics engine regarded as a black box.
@@
    GASP does not employ additional meshing strategies. We demonstrate that applying additional Gaussian control rules effectively eliminates artifacts, while leveraging hierarchy significantly accelerates simulations.
@@ 2 Related Work
The GS representation of 3D scenes is suitable for editing [5, 8, 22] and modeling dynamic scenes [11, 23, 21]. However, the more challenging task is its integration with a physics engine that remains difficult and underexplored.
Figure 3: GASP produces physical simulations using flat GS representation. It works with static GS models like GaMeS [22] or dynamic ones like D-MiSo [21]. First, each flat Gaussian component is converted to three points (a triangle face). Then, the physics engine is run on a selected point cloud to obtain trajectories for our 3D model. Finally, we convert points into Gaussian components by reverse GaMeS parametrization. Notably, we also provided some additional rules that can be applied to control Gaussians during simulations.
@@
Most existing methods use additional strategic tactics to obtain representations better aligned with physics engines’ requirements. Gaussian Splashing (GSP) [4] is a unified framework that combines 3D Gaussian Splatting (3DGS) and position-based dynamics. During the preprocessing, the foreground objects are isolated and reconstructed as a mesh. Such a response is then combined with position-based dynamics. Case-specific methods, such as Embodied Gaussians [1], introduce a Gaussian Particle that represents an object through particles in order to maintain its geometric aspects for robotics applications. Alternatively, we can use cage-based representation [12]. Other methods, such as LITA-GS [26], extract physical attributes from scenes in order to adapt the lighting conditions. VR-GS [12] combines 3DGS and eXtended Position-based Dynamics (XPBD) [16]. The model starts with scene reconstruction by GS and tetrahedralization. Consequently, we obtain a cage-based representation merged with a physics engine.
@@
All of the above models work with an additional meshing strategy. In contrast, PhysGaussian [24] operates directly on Gaussian distributions. The authors of [24] suggest approximating a deformation map using the first-order derivative. This yields a linear transformation capable of modifying Gaussian components. Although this model produces relatively good results, it requires adjustments to the physics engine. In contrast, we propose GASP model, which uses a physics engine as a black box without any modification.
@@
A separate area in physics modeling involves using diffusion models to generate object dynamics. PhysDreamer [25] introduces a method that provides static 3D objects with interactive dynamics by utilizing object dynamics priors acquired from video generation models. Physics3D [15] learns various physical characteristics of 3D objects using a video diffusion model. DreamPhysics [10] infers the physical attributes of 3DGS based on video diffusion priors.
3 GASP pipeline Overview
@@
Here, we present a description of the GASP pipeline (see Fig. 3). It begins with a general GS model, followed by an examination of the modifications made by GaMeS [22] to the original GS model, which facilitates straightforward adaptations. Subsequently, the classical MPM [6] is outlined, and finally, our own approach combining both GS and MPM is detailed. In order to improve visual and time quality, we have also proposed two strategies. One is based on the control of the size of Gaussians, and the other is on making appropriate Core Gaussians to create hierarchies (see Fig. 4).
@@
Figure 4: In the GASP pipeline, the hierarchical organization of Gaussians through clustering methods allows simulations to focus only on Gaussians at the highest levels of the hierarchy. This approach speeds up computations, especially when using particle-based engines such as Taichi.
In the classical GS model, each element is characterized by a collection of parameters, including a covariance matrix $\Sigma$ which is factored as $\Sigma=RSS^{T}R^{T}$ where $R$ is the rotation matrix and $S$ is a diagonal matrix containing the scaling parameters. In GaMeS [22], on the other hand, flat Guassians are used $(\mathcal{N}({\bf m},R,S),\sigma,c),$ where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$, with $s_{1}=\varepsilon$, and $R$ is the rotation matrix defined as $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$, with ${\bf r}_{i}\in\mathbb{R}^{3}$. Thus, Gaussian components can be visualized through a triangle-face mesh by utilizing the vertices for their parameterization. We refer to this mapping as $\mathcal{T}(\cdot)$. In practice, this parameterization results in a set of triangles, which is referred to as triangle soup [18].
To outline the GaMeS parameterization, consider a single Gaussian component $\mathcal{N}({\bf m},R,S)$. Then its face representation $\mathcal{N}(V)$ is based on a triangle $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]=\mathcal{T}({\bf m},R,S)$ with the vertices defined as ${\bf v}_{1}={\bf m},\;{\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2},\;\textrm{and}\;{\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}.$ Conversely, given a face (triangle) representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$, we can recover the Gaussian component:
@@
Here, $\mathrm{orth}(\cdot)$ denotes a single step of the Gram-Schmidt process [3]. Accordingly, the corresponding covariance matrix of a Gaussian distribution is given as $\hat{\Sigma}=\hat{R}\hat{S}\hat{S}^{T}\hat{R}^{T}.$
In addition, GaMeS supports distributing Gaussians on meshes when a mesh structure is provided. In this setting, Gaussian means are assigned to each mesh face via learnable barycentric coordinates of its vertices, ensuring they remain bound to the mesh under deformations. The orientation of each Gaussian is defined by the face normal and two in-plane directions derived from the vertices, allowing consistent updates when the mesh position changes. In this work, we focus on the mesh-free case, though the proposed formulations are directly applicable to Gaussians placed on mesh faces.
$t_{0}$ $t_{i}$ $t_{j}$ $t_{T}$
2DGS Refer to caption Refer to caption Refer to caption Refer to caption
GaMeS Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Visual comparison of gravity simulations with sand-like material represented by points. First row: direct application of physical simulators to Gaussians obtained from 2DGS, which produces noticeable artifacts. Second row: application of physical simulators on triangle faces obtained using the GaMeS parametrization, the pointed artifacts and holes are not as apparent.
@@
, which is used to measure local rotation and strain. The deformation function $\phi$ is subject to changes in accordance with the principles of conservation of mass, conservation of momentum, and the elasto-plastic constitutive relation:
@@
Figure 6: GASP operates directly on the Gaussian triangle representation without altering the physics-based simulation engine. Initially, the 3D objects are depicted as flat Gaussians. GaMeS then converts these Gaussians into triplets (mesh faces). The simulation engine transforms these points following the MPM model. Subsequently, the new points are reassembled into triangles and then into Gaussians. This process is speedy, yielding realistic renderings.
@@
This subsection presents a description of the GASP pipeline. Since the model employs the GaMeS representation of a 3D scene, we postulate that the input comprises a collection of Gaussians:
@@
with means ${\bf m}_{i}$, scaling matrices $S_{i}$, and rotation matrices $R_{i}$. Furthermore, we assume the existence of a deformation map $\phi(X,t)$, which models the behavior of 3D objects. To obtain $\phi(X,t)$, it is necessary to have a point cloud that describes 3D objects. As our method employs the GaMeS reparametrization of GS, we can transform the mean Gaussian parameters and the covariance matrix into a mesh (triangle soup). In such an instance, we may consider the system $(\mathcal{V},\mathcal{E})$, where:
@@
denotes a collection of vertices, and $\mathcal{E}$ represents the edges obtained via GaMeS. Our physics-based simulation engine operates at each point of $V_{i}$ individually. Subsequently, the vertices of the triangle soup can be used as input for the MPM algorithm in conjunction with tools such as the Blender software, Taichi_elements, or Genesis libraries (for further details, see the experimental section), see Fig. 2. It is essential to note that our solution does not entail any modifications to the MPM, in contrast to the Taylor approximation employed in [24]. Consequently, in each time step, the following is obtained:
@@
The triangle soup in time $t$ is then defined by the collection of vertices and edges $(\mathcal{V}_{t},\mathcal{E}_{t})$. In contrast, the reverse parametrization provided by GaMeS can be employed to obtain Gaussian Splatting:
@@
Our GASP pipeline represents Gaussian components through triangle faces and integrates them with a point-based physics engine. Flat Gaussian components are first converted into triangle faces, and the simulation is then performed on their vertices. Direct application of standard physics simulators often produces noticeable artifacts (see Fig. 5). In contrast, our approach effectively supports the learned flat Gaussians by transforming them into triangles, ensuring stable simulations. Unlike alternative flat Gaussian representations, such as 2DGS, the GaMeS parametrization naturally fits this setting, since each flat Gaussian is inherently represented by a triangle.
Figure 7: Comparison of physical animation models, illustrating the effects of applying an algorithm to correct points used in simulation. The left side demonstrates a simulation without the correction algorithm, while the right side shows the enhanced accuracy and stability achieved using Taichi elements with the correction algorithm applied.
Figure 8: GASP works directly on points cloud extracted from flat Gaussians. In practice, physics engine like Taichi_elements moves all particles independently, which can produce artifacts. Therefore, we use a correction algorithm to control the distance of points from each other. If some points are moved extremely far, we automatically correct such behaviors by clipping parameters $S$ in GS.
Gaussian Control Rules
It is important to note that processing each vertex of a triangle soup individually is a highly rapid process that treats the physics engine as a black box. The schema of such processes is presented in Fig. 6. Our method is compatible with any physics engine that operates on three-dimensional points. However, the processing of all points separately may yield artifacts. When all points are treated independently during simulation, some elements exhibit considerable fluctuations in velocity. This phenomenon leads to the formation of extensive triangles (mesh faces), which in turn give rise to prominent Gaussian components, as illustrated in Fig. 7. To mitigate this effect, our approach employs a straightforward yet highly effective technique (correction algorithm). Specifically, we monitor the distance between vertices of the face representation $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ of a Gaussian component $\mathcal{N}({\bf m},R,S)$. It should be noticed that, in the case of GaMeS transformation, the first vertex is equal to the center of the distribution (${\bf v}_{1}={\bf m}$), while the two subsequent vertices are located on the main components of the ellipses that describe the distribution (${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$). Consequently, if the distance between the transformed mesh nodes is considerably greater than that of the original nodes, i.e.:
@@
where $\alpha$ is a hyperparameter, we automatically assert that the value of $s_{i}$ should be identical to that of the original value times $\alpha$. This is illustrated in Fig. 8.
Gaussians Hierarchy in Simulations
Figure 9: Blender provides a wide range of physical properties and force sources, including gravity and wind, to create realistic and dynamic animations. In this experiment, we present examples of physics animation in object scenes, showing the versatility of animating both small objects, such as a flower on a table, and large objects, such as a fox’s head dominating the scene.
Achieving high-quality renders with Gaussian Splatting methods typically requires a substantial number of Gaussians. Previously, animations were executed directly on Gaussians [24] or using estimated meshes [5]. However, these simulations demand significant time and computational resources. In the Genesis engine [2], this process is optimized by employing simulations facilitated through particles selected via Sampler component, lowering the number of points. To achieve a similar mechanism, we employ BIRCH clustering to construct Gaussians called Core-Gaussians for each cluster. This enables the creation of a hierarchical structure for Gaussians. The simulation is then performed only on Core Gaussians, and then their move pattern is transferred on Sub Gaussians inspired by [21], see Fig. 4. We demonstrate that physics engines can be applied to selected Gaussians without compromising visual quality (see Fig. 4). This approach significantly accelerates the simulation process while maintaining rendering quality.
@@
GASP is a natural combination of a physics engine dedicated to points and Gaussian Splatting with flat Gaussians. The reparametrization in GaMeS can be applied to a multitude of disparate methodologies, with a particular focus on flat Gaussian methods. Moreover, it can be employed for dynamic Gaussian Splatting. In [21], the D-MiSo model is introduced, which uses flat Gaussians for the modeling of dynamic scenes. Such representations can be utilized in conjunction with the aforementioned approach. In the D-MiSo framework, two distinct types of Gaussians and two deformable networks are employed. In the process of inference, only sub-Gaussians are employed, which are modified by deformable function composition $\psi(v,s)$, i.e.,
@@
where $s$ represents a time value derived from the dynamics observed in the scenes. In our solution, we can incorporate a deformation map $\phi(X,t)$, which comes from the physics simulator. Consequently, GASP for the dynamic scene is given as follows: